# LASER: Layer-wise Scale Alignment for Training-Free Streaming 4D Reconstruction

### v1

@@ 1 Introduction
Recovering 3D scene geometry from images has long been a central pursuit in computer vision, with applications ranging from robotic perception to digital cultural preservation. For decades, this problem was addressed through geometry-centric pipelines: Structure-from-Motion (SfM) [hartley2003multiple, schoenberger2016sfm] and Multi-View Stereo (MVS) [furukawa2009accurate, schoenberger2016mvs] systems with hand-crated designs. While these classical methods achieve impressive accuracy with careful engineering, they remain sensitive to textureless regions and require known or estimated camera calibration. Recent advent of feed-forward neural approaches have fundamentally changed this landscape. DUSt3R [dust3r] pioneers direct regression of dense pointmaps from uncalibrated image pairs, eliminating the need for explicit correspondence solving. Subsequent work including VGGT [wang2025vggt] and $\pi^{3}$ [wang2025pi] further handle arbitrary numbers of views, establishing a new paradigm where large-scale transformers trained on diverse data achieve superior reconstruction quality in zero-shot settings.
@@
While these offline models achieve excellent reconstruction quality, they struggle with streaming scenarios due to quadratic memory complexity and the need to reprocess all frames when new observations arrive. Several recent works have proposed streaming variants that process frames incrementally. Some approaches [wang2025cut3r, wang20243d, chen2025long3r] introduce persistent state or memory mechanisms for continuous 3D perception. Another line of works [zhuo2025streaming, stream3r2025, li2025wint3r] adapt offline models with causal attention or combine sliding window with camera token pools. Though effective, these methods share a common limitation: they require extensive retraining from scratch or through knowledge distillation to learn streaming-setting reconstruction, which is computationally expensive and may not fully leverage the strong geometric priors of state-of-the-art offline models like VGGT [wang2025vggt] and $\pi^{3}$ [wang2025pi]. Moreover, recurrent designs like CUT3R [wang2025cut3r] can suffer from drift and catastrophic forgetting over long sequences [chen2025ttt3r], while methods relying on growing memory face scalability constraints. Concurrent work VGGT-Long [deng2025vggtlongchunkitloop] also pursues a training-free approach by chunking sequences and aligning with $Sim(3)$, but as we show in Sec. 4, simple rigid alignment is insufficient. Given that offline models already encode rich 3D priors, we ask: can we achieve both training-free conversion for streaming input and robust geometric alignment?
@@
In this work, we propose LASER, a training-free framework that converts offline models into streaming systems by revisiting classical geometric principles. LASER  employs a sliding-window strategy, processing overlapping subsets of frames (windows) sequentially with a frozen offline model as the backbone. Modern feed-forward models like VGGT [wang2025vggt] and $\pi^{3}$ [wang2025pi] excel at producing accurate 3D reconstructions within individual windows. However, aligning these windows consistently remains challenging. We observe that simple $Sim(3)$ alignment fails due to layer depth misalignment: monocular scale ambiguity causes relative depth scales across scene layers (e.g., foreground vs. background) to vary between windows, particularly when camera translation is limited. A global $Sim(3)$ transformation applies uniform scaling to the entire window and cannot resolve such layer-wise scale variations. Drawing on classical insights that scenes naturally decompose into depth-ordered layers with distinct geometric properties [shade1998layered, wang1994representing, baker1998layered], we propose layer-wise scale alignment adapted to the modern deep learning reconstruction setting. Our approach segments reconstructed point maps into discrete layers using [felzenszwalb2004efficient], computes per-layer scale factors between consecutive windows, and propagates these scales across the sequence to achieve layer-consistent alignment.
@@
Experimental results show that our design effectively addresses the aforementioned challenges and achieves state-of-the-art performance. LASER  outperforms the learned streaming methods while processing image streams at 14 FPS with only 6 GB peak memory on one RTX A6000 GPU. Notably, our training-free approach maintains competitive reconstruction quality with offline models (0.013mm vs 0.011mm mean accuracy on 7-Scenes [seven_scenes]) while enabling online processing. This shows that when deep learning models provide strong local geometry, classical layer-based geometric reasoning can effectively unify their outputs into consistent long-range reconstructions without retraining. Beyond quantitative gains, LASER offers significant practical advantages: it requires no model retraining, immediately applies to existing offline reconstruction models (as demonstrated with both VGGT [wang2025vggt] and $\pi^{3}$ [wang2025pi] backbones), and scales to kilometer-long sequences that exceed the memory capacity of offline methods. As new, more powerful offline models emerge, LASER can immediately leverage their improvements without additional training costs, bridging the gap between the quality and efficiency of offline models and the streaming requirements.
@@
  • We propose LASER, a training-free framework for converting offline reconstruction models into streaming systems without retraining, (e.g., VGGT [wang2025vggt], $\pi^{3}$ [wang2025pi]).
@@
  • LASER achieves state-of-the-art performance in streaming pose estimation and 3D reconstruction ( $68.6\%$ ATE on Sintel [sintel] pose estimation, $63.9\%$ Acc on 7-Scenes [seven_scenes] reconstruction compared to the previous best) while operating at 14 FPS on an A6000 GPU with only 6 GB peak runtime memory.
@@ 2 Related Work
Learning-based 3D Reconstruction. Learning-based methods recast 3D reconstruction as a data-driven estimation problem rather than a purely geometric one. Early CNN-based pipelines [ji2017surfacenet, DeepMVS, yao2018mvsnet] replace handcrafted correspondence matching with learned feature aggregation and differentiable depth regression, paving the way toward end-to-end multi-view geometry learning. Subsequent methods [murez2020atlas, sun2021neucon] extend these ideas to online or large-scale reconstruction via recurrent fusion and volumetric integration. Implicit-field formulations [mildenhall2020nerf, wang2021neus, yu2022monosdf] achieve photorealistic surface and appearance modeling from sparse or monocular inputs, while explicit representations [kerbl3Dgaussians, triplane] further improve rendering efficiency and scalability.However, these models typically require per-scene optimization and are bounded by scene complexities [liang2025monocular]. A parallel line of research seeks to eliminate the optimization through feed-forward geometric reasoning. DUSt3R [dust3r] pioneers a paradigm where 3D point clouds and relative poses are directly regressed from image pairs, and VGGT [wang2025vggt] generalizes this idea to variable-sized image sets with the help of learnable camera tokens. The recent $\pi^{3}$ model [wang2025pi] further introduces permutation-equivariant attention to unify structure and motion within a single scalable framework. Our proposed LASER builds on this feed-forward foundation, introducing a lightweight streaming formulation that adapts offline reconstruction models for continuous, efficient, kilometer-scale processing without retraining.
Learning-based 4D Reconstruction. Learning-based 4D reconstruction approaches extend geometric and appearance modeling to temporal domain. Early progress built on implicit neural representations [pumarola2020d, park2021nerfies, park2021hypernerf, li2020neural] introduces temporal conditioning or deformation fields. To improve efficiency, subsequent approaches [WU_2024_CVPR, yang2023deformable3dgs, luiten2023dynamic, gaufre, kplanes_2023, Cao2023HexPlane, TiNeuVox, liu2024gear] apply similar extensions to explicit representations. Both implicit and explicit 4D representations require per-scene optimization , limiting usage in streaming settings. Another series of works [han2025enhancing, zhang2025monstr, lu2024align3r, chen2025easi3r] extend feed-forward regression to dynamic settings. [st4rtrack2025, jin2024stereo4d, Liang2025ZeroShotMSF, li2024_MegaSaM] leverage dense video correspondence supervision to generalize to diverse dynamic scenes. $\pi^{3}$ [wang2025pi] unifies 3D and 4D reasoning through permutation-equivariant attention during large-scale training. Our method shares the goal of scalable 4D reconstruction but focuses on the streaming regime: adapting offline feed-forward geometry transformers for causal video processing. With a sliding-window formulation and geometry-aware alignment, LASER achieves temporally consistent reconstruction efficiently without retraining.
@@
Streaming Feed-Forward Reconstruction. Real-world applications such as autonomous driving, robotics, and AR/VR require models that process video streams efficiently and consistently. Recent works have tried to finetune offline feed-forward reconstructor to the streaming regime: memory-centric approaches introduce persistent or explicit spatial memory to extend temporal horizons [wang2025cut3r, wu_point3r_2025, spann3r]; causal-transformer designs process frames sequentially with token pooling or causal attention [li2025wint3r, stream3r2025, zhuo2025streaming]; test-time adaptation has also been explored for long videos [chen2025ttt3r]; parallel efforts bridge feed-forward prediction with SLAM-style optimization [mast3r_slam, vggtslam]. Aside from requiring retraining, these methods either accumulate scale drift over time, fail on long sequences due to memory limits or incur slow inference. On the other hand, [Yang_2025_Fast3R] pushes feed-forward reconstruction to the kilo-frame regime, but does not support streaming input. LASER addresses these limitations with a general, training-free framework that turns offline feed-forward models (e.g., VGGT or $\pi^{3}$) into a streaming system capable of handling kilo-frame dynamic sequences while preserving their reconstruction quality and efficiency.
@@ 3.1 Overview
For each window $\mathcal{W}_{i}$, a pretrained feed-forward reconstructor $f(\cdot)$ (e.g., VGGT [wang2025vggt], $\pi^{3}$ [wang2025pi]) predicts per-frame point maps and camera poses:
@@
To estimate the $\mathrm{Sim}(3)$ transform, we first estimate the global scale factor $s_{i}^{w}$ via a robust IRLS (Iteratively Reweighted Least Squares) optimization [irls], enforcing a shared metric across two adjacent windows. Rotation and translation $(\mathbf{R}_{i}^{w},\mathbf{t}_{i}^{w})$ are then optimized via the Kabsch algorithm  [Kabsch] under that metric using the scaled camera anchors based on the estimated $s_{i}^{w}$. We refer readers to the supplementary material for more details.
@@ 3.2 Layer-wise Scale Alignment (LSA)
Although the global $\mathrm{Sim}(3)$ registration aligns each window to a common scale, it assumes isotropic scaling, where the same scale factor applies equally along all spatial axes. In practice, this assumption breaks, e.g., under low-parallax motion, where a monocular reconstructor cannot reliably constrain depth (the $Z$-axis) relative to lateral axes. As a result, even after the global alignment, surfaces at different depths may exhibit layer-wise scale inconsistency: foreground regions appear over- or under-scaled relative to background structures across windows, as shown in Fig. 3. This anisotropic scaling along depth accumulates over time, leading to visible distortions and metric drift in the fused reconstruction. Following classical insights that scenes decompose into depth-ordered layers [shade1998layered, baker1998layered], we introduce Layer-wise Scale Alignment (LSA), a geometry-driven refinement that corrects distortions based on a layer graph.
Depth layer extraction. Inspired by classical layered representations, where a scene is decomposed into depth-ordered surfaces [shade1998layered, baker1998layered], we extract depth layers by segmenting each depth map into spatially coherent regions at similar depths. Specifically, let $\bar{\mathbf{P}}_{t}^{(i)}=\mathbf{T}_{t}^{w}\mathbf{P}_{t}^{(i)}\!\in\!\mathbb{R}^{H\times W\times 3}$ denote the 3D point map after the $\mathrm{Sim}(3)$ registration for frame $\mathbf{I}_{t}$ in the temporal window $\mathcal{W}_{i}$. We derive a pseudo-depth map, denoted as $\bar{\mathbf{D}}_{t}^{(i)}$, from $\bar{\mathbf{P}}_{t}^{(i)}$ by taking its $Z$-coordinate components. This pseudo-depth map is partitioned into $M_{t}^{(i)}$ disjoint depth layers $\{\mathcal{L}_{t,m}^{(i)}\}_{m=1}^{M_{t}^{(i)}}$ using an efficient segmentation algorithm [felzenszwalb2004efficient]. Each layer $\mathcal{L}_{t,m}^{(i)}$ corresponds to a continuous geometric surface patch with a coherent depth. Examples are shown in Fig. 3.
@@ 4.1.1 Tasks, datasets, and metrics
Following [zhang2025monstr, wang2025cut3r], we evaluate on Sintel [sintel], Bonn [bonn], and KITTI [kitti]. Predicted depths are aligned to ground truth via scale-only alignment.
Camera pose estimation protocol. Following [zhang2025monstr, wang2025cut3r], we compare on small-scale Sintel [sintel], ScanNet [scannet], and TUM RGB-D [tum_rgbd]. Predicted trajectories are aligned to ground truth via a Sim(3) transformation. For large-scale evaluation, we use KITTI Odometry [kitti] following [deng2025vggtlongchunkitloop].
Multi-view point map estimation protocol. We run evaluation on 7-Scenes [seven_scenes] and NRGBD [nrgbd] with a keyframe sampling interval of 10, except for using an interval of 15 on NRGBD for [zhuo2025streaming, stream3r2025] due to their memory limits. Predicted point maps are registered to ground truth using the Umeyama algorithm (coarse Sim(3) alignment) followed by Iterative Closest Point (ICP) refinement.
Table 1: Video Depth Estimation on Sintel [sintel], Bonn [bonn] and KITTI [kitti]. We report Abs Rel and $\delta{<}1.25$.
@@
VGGT [wang2025vggt] 0.303 68.5 0.055 97.1 0.073 96.3
$\pi^{3}$ [wang2025pi] 0.245 68.4 0.050 97.5 0.038 98.6
Spann3R [spann3r] 0.622 42.6 0.144 81.3 0.198 73.7
CUT3R [wang2025cut3r] 0.421 47.9 0.078 93.7 0.118 88.1
Point3R [wu_point3r_2025] 0.452 48.9 0.060 96.0 0.136 84.2
VGGT-SLAM [vggtslam] 0.424 56.0 0.076 93.2 0.136 81.8
StreamVGGT [zhuo2025streaming] 0.323 65.7 0.059 97.2 0.173 72.1
@@
Table 2: Camera Pose Estimation on Sintel [sintel], ScanNet [scannet], and TUM [tum_rgbd]. We report ATE, translational RPE, and rotational RPE.
@@
DUSt3R [dust3r] 0.290 0.132 7.869 0.246 0.108 8.210 0.140 0.106 3.286
VGGT [wang2025vggt] 0.171 0.062 0.471 0.035 0.015 0.381 0.012 0.010 0.309
$\pi^{3}$ [wang2025pi] 0.073 0.037 0.287 0.030 0.012 0.346 0.014 0.009 0.307
Spann3R [spann3r] 0.329 0.110 4.471 0.096 0.023 0.661 0.056 0.021 0.591
CUT3R [wang2025cut3r] 0.213 0.066 0.621 0.099 0.022 0.600 0.046 0.015 0.473
Point3R [wu_point3r_2025] 0.351 0.128 1.822 0.106 0.035 1.946 0.075 0.029 0.642
VGGT-SLAM [vggtslam] 0.303 0.128 6.883 0.070 0.049 2.447 0.030 0.020 1.567
StreamVGGT [zhuo2025streaming] 0.251 0.149 1.894 0.161 0.057 3.647 0.061 0.033 3.209
@@
Table 3: Large-scale Camera Pose Estimation on KITTI [kitti]. We report ATE (lower is better). CF: checkmark () indicates no calibration required; DR: checkmark () indicates dense reconstruction supported. We show metrics for sequence ID; Avg. is the mean across sequences. Seq. 01 corresponds to a high-speed driving sequence whose motion differs from others; Avg. reports the mean ATE excluding Seq. 01. OOM: CUDA out-of-memory, TL: tracking lost.
@@
ORB-SLAM2 (w/o LC) [murORB2] 69.73 26.48 40.65 502.20 47.82 0.94 1.30 29.95 40.82 16.04 43.09 38.77 5.42
ORB-SLAM2 (w/ LC) [murORB2] 54.82 9.46 6.03 508.34 14.76 1.02 1.57 4.04 11.16 2.19 38.85 8.39 6.63
LDSO [gao2018ldso] 22.43 23.50 9.32 11.68 31.98 2.85 1.22 5.10 13.55 2.96 129.02 21.64 17.36
DROID-VO [teed2021droid] 54.19 51.19 98.43 84.20 108.80 2.58 0.93 59.27 64.40 24.20 64.55 71.80 16.91
DPVO [teed2023deep] 53.61 57.70 113.21 12.69 123.40 2.09 0.68 58.96 54.78 19.26 115.90 75.10 13.63
DROID-SLAM [teed2021droid] 100.28 75.85 92.10 344.60 107.61 2.38 1.00 118.50 62.47 21.78 161.60 72.32 118.70
DPV-SLAM [lipson2024deep] 53.03 57.19 112.80 11.50 123.53 2.50 0.81 57.80 54.86 18.77 110.49 76.66 13.65
DPV-SLAM++ [lipson2024deep] 25.75 27.14 8.30 11.86 39.64 2.50 0.78 5.74 11.60 1.52 110.90 76.70 13.70
VGGT [wang2025vggt] / / OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
$\pi^{3}$ [wang2025pi] / / OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
MASt3R-SLAM [mast3r_slam] / / TL TL TL TL TL TL TL TL TL TL TL
CUT3R [wang2025cut3r] / / OOM OOM OOM 148.07 22.31 OOM OOM OOM OOM OOM OOM
Fast3R [Yang_2025_Fast3R] / / OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
VGGT-Long [deng2025vggtlongchunkitloop] 27.64 18.28 8.67 121.17 32.08 6.12 4.23 8.31 5.34 4.63 53.10 41.99 18.37
@@
Table 5: Ablation Studies for Layer-wise Scale Alignment (Sec. 3.2). wo// LSA denotes not using LSA component. w// SAM 2 denotes replacing the efficient segmentation algorithm [felzenszwalb2004efficient] with a recent counterpart SAM 2. wo// $\mathcal{E}_{\mathrm{intra}}$ denotes removing the temporal propagation step from LSA.

### v2

@@ 1 Introduction
Recovering 3D scene geometry from images has long been a central pursuit in computer vision, with applications ranging from robotic perception to digital cultural preservation. For decades, this problem was addressed through geometry-centric pipelines: Structure-from-Motion (SfM) [18, 48] and Multi-View Stereo (MVS) [14, 49] systems with hand-crated designs. While these classical methods achieve impressive accuracy with careful engineering, they remain sensitive to textureless regions and require known or estimated camera calibration. Recent advent of feed-forward neural approaches have fundamentally changed this landscape. DUSt3R [63] pioneers direct regression of dense pointmaps from uncalibrated image pairs, eliminating the need for explicit correspondence solving. Subsequent work including VGGT [59] and $\pi^{3}$ [64] further handle arbitrary numbers of views, establishing a new paradigm where large-scale transformers trained on diverse data achieve superior reconstruction quality in zero-shot settings.
@@
While these offline models achieve excellent reconstruction quality, they struggle with streaming scenarios due to quadratic memory complexity and the need to reprocess all frames when new observations arrive. Several recent works have proposed streaming variants that process frames incrementally. Some approaches [62, 57, 7] introduce persistent state or memory mechanisms for continuous 3D perception. Another line of works [73, 24, 27] adapt offline models with causal attention or combine sliding window with camera token pools. Though effective, these methods share a common limitation: they require extensive retraining from scratch or through knowledge distillation to learn streaming-setting reconstruction, which is computationally expensive and may not fully leverage the strong geometric priors of state-of-the-art offline models like VGGT [59] and $\pi^{3}$ [64]. Moreover, recurrent designs like CUT3R [62] can suffer from drift and catastrophic forgetting over long sequences [6], while methods relying on growing memory face scalability constraints. Concurrent work VGGT-Long [9] also pursues a training-free approach by chunking sequences and aligning with $Sim(3)$, but as we show in Sec. 4, simple rigid alignment is insufficient. Given that offline models already encode rich 3D priors, we ask: can we achieve both training-free conversion for streaming input and robust geometric alignment?
@@
In this work, we propose LASER, a training-free framework that converts offline models into streaming systems by revisiting classical geometric principles. LASER  employs a sliding-window strategy, processing overlapping subsets of frames (windows) sequentially with a frozen offline model as the backbone. Modern feed-forward models like VGGT [59] and $\pi^{3}$ [64] excel at producing accurate 3D reconstructions within individual windows. However, aligning these windows consistently remains challenging. We observe that simple $Sim(3)$ alignment fails due to layer depth misalignment: monocular scale ambiguity causes relative depth scales across scene layers (e.g., foreground vs. background) to vary between windows, particularly when camera translation is limited. A global $Sim(3)$ transformation applies uniform scaling to the entire window and cannot resolve such layer-wise scale variations. Drawing on classical insights that scenes naturally decompose into depth-ordered layers with distinct geometric properties [50, 60, 1], we propose layer-wise scale alignment adapted to the modern deep learning reconstruction setting. Our approach segments reconstructed point maps into discrete layers using [11], computes per-layer scale factors between consecutive windows, and propagates these scales across the sequence to achieve layer-consistent alignment.
@@
Experimental results show that our design effectively addresses the aforementioned challenges and achieves state-of-the-art performance. LASER  outperforms the learned streaming methods while processing image streams at 14 FPS with only 6 GB peak memory on one RTX A6000 GPU. Notably, our training-free approach maintains competitive reconstruction quality with offline models (0.013mm vs 0.011mm mean accuracy on 7-Scenes [51]) while enabling online processing. This shows that when deep learning models provide strong local geometry, classical layer-based geometric reasoning can effectively unify their outputs into consistent long-range reconstructions without retraining. Beyond quantitative gains, LASER offers significant practical advantages: it requires no model retraining, immediately applies to existing offline reconstruction models (as demonstrated with both VGGT [59] and $\pi^{3}$ [64] backbones), and scales to kilometer-long sequences that exceed the memory capacity of offline methods. As new, more powerful offline models emerge, LASER can immediately leverage their improvements without additional training costs, bridging the gap between the quality and efficiency of offline models and the streaming requirements.
@@
  • We propose LASER, a training-free framework for converting offline reconstruction models into streaming systems without retraining, (e.g., VGGT [59], $\pi^{3}$ [64]).
@@
  • LASER achieves state-of-the-art performance in streaming pose estimation and 3D reconstruction ( $68.6\%$ ATE on Sintel [2] pose estimation, $63.9\%$ Acc on 7-Scenes [51] reconstruction compared to the previous best) while operating at 14 FPS on an A6000 GPU with only 6 GB peak runtime memory.
@@ 2 Related Work
Learning-based 3D Reconstruction. Learning-based methods recast 3D reconstruction as a data-driven estimation problem rather than a purely geometric one. Early CNN-based pipelines [20, 19, 70] replace handcrafted correspondence matching with learned feature aggregation and differentiable depth regression, paving the way toward end-to-end multi-view geometry learning. Subsequent methods [40, 53] extend these ideas to online or large-scale reconstruction via recurrent fusion and volumetric integration. Implicit-field formulations [37, 61, 71] achieve photorealistic surface and appearance modeling from sparse or monocular inputs, while explicit representations [23, 4] further improve rendering efficiency and scalability.However, these models typically require per-scene optimization and are bounded by scene complexities [30]. A parallel line of research seeks to eliminate the optimization through feed-forward geometric reasoning. DUSt3R [63] pioneers a paradigm where 3D point clouds and relative poses are directly regressed from image pairs, and VGGT [59] generalizes this idea to variable-sized image sets with the help of learnable camera tokens. The recent $\pi^{3}$ model [64] further introduces permutation-equivariant attention to unify structure and motion within a single scalable framework. Our proposed LASER builds on this feed-forward foundation, introducing a lightweight streaming formulation that adapts offline reconstruction models for continuous, efficient, kilometer-scale processing without retraining.
Learning-based 4D Reconstruction. Learning-based 4D reconstruction approaches extend geometric and appearance modeling to temporal domain. Early progress built on implicit neural representations [46, 43, 44, 25] introduces temporal conditioning or deformation fields. To improve efficiency, subsequent approaches [65, 69, 34, 29, 13, 3, 10, 32] apply similar extensions to explicit representations. Both implicit and explicit 4D representations require per-scene optimization , limiting usage in streaming settings. Another series of works [17, 72, 33, 5] extend feed-forward regression to dynamic settings. [12, 21, 28, 26] leverage dense video correspondence supervision to generalize to diverse dynamic scenes. $\pi^{3}$ [64] unifies 3D and 4D reasoning through permutation-equivariant attention during large-scale training. Our method shares the goal of scalable 4D reconstruction but focuses on the streaming regime: adapting offline feed-forward geometry transformers for causal video processing. With a sliding-window formulation and geometry-aware alignment, LASER achieves temporally consistent reconstruction efficiently without retraining.
@@
Streaming Feed-Forward Reconstruction. Real-world applications such as autonomous driving, robotics, and AR/VR require models that process video streams efficiently and consistently. Recent works have tried to finetune offline feed-forward reconstructor to the streaming regime: memory-centric approaches introduce persistent or explicit spatial memory to extend temporal horizons [62, 66, 58]; causal-transformer designs process frames sequentially with token pooling or causal attention [27, 24, 73]; test-time adaptation has also been explored for long videos [6]; parallel efforts bridge feed-forward prediction with SLAM-style optimization [39, 35]. Aside from requiring retraining, these methods either accumulate scale drift over time, fail on long sequences due to memory limits or incur slow inference. On the other hand, [68] pushes feed-forward reconstruction to the kilo-frame regime, but does not support streaming input. LASER addresses these limitations with a general, training-free framework that turns offline feed-forward models (e.g., VGGT or $\pi^{3}$) into a streaming system capable of handling kilo-frame dynamic sequences while preserving their reconstruction quality and efficiency.
@@ 3.1 Overview
For each window $\mathcal{W}_{i}$, a pretrained feed-forward reconstructor $f(\cdot)$ (e.g., VGGT [59], $\pi^{3}$ [64]) predicts per-frame point maps and camera poses:
@@
To estimate the $\mathrm{Sim}(3)$ transform, we first estimate the global scale factor $s_{i}^{w}$ via a robust IRLS (Iteratively Reweighted Least Squares) optimization [41], enforcing a shared metric across two adjacent windows. Rotation and translation $(\mathbf{R}_{i}^{w},\mathbf{t}_{i}^{w})$ are then optimized via the Kabsch algorithm  [22] under that metric using the scaled camera anchors based on the estimated $s_{i}^{w}$. We refer readers to the supplementary material for more details.
@@ 3.2 Layer-wise Scale Alignment (LSA)
Although the global $\mathrm{Sim}(3)$ registration aligns each window to a common scale, it assumes isotropic scaling, where the same scale factor applies equally along all spatial axes. In practice, this assumption breaks, e.g., under low-parallax motion, where a monocular reconstructor cannot reliably constrain depth (the $Z$-axis) relative to lateral axes. As a result, even after the global alignment, surfaces at different depths may exhibit layer-wise scale inconsistency: foreground regions appear over- or under-scaled relative to background structures across windows, as shown in Fig. 3. This anisotropic scaling along depth accumulates over time, leading to visible distortions and metric drift in the fused reconstruction. Following classical insights that scenes decompose into depth-ordered layers [50, 1], we introduce Layer-wise Scale Alignment (LSA), a geometry-driven refinement that corrects distortions based on a layer graph.
Depth layer extraction. Inspired by classical layered representations, where a scene is decomposed into depth-ordered surfaces [50, 1], we extract depth layers by segmenting each depth map into spatially coherent regions at similar depths. Specifically, let $\bar{\mathbf{P}}_{t}^{(i)}=\mathbf{T}_{t}^{w}\mathbf{P}_{t}^{(i)}\!\in\!\mathbb{R}^{H\times W\times 3}$ denote the 3D point map after the $\mathrm{Sim}(3)$ registration for frame $\mathbf{I}_{t}$ in the temporal window $\mathcal{W}_{i}$. We derive a pseudo-depth map, denoted as $\bar{\mathbf{D}}_{t}^{(i)}$, from $\bar{\mathbf{P}}_{t}^{(i)}$ by taking its $Z$-coordinate components. This pseudo-depth map is partitioned into $M_{t}^{(i)}$ disjoint depth layers $\{\mathcal{L}_{t,m}^{(i)}\}_{m=1}^{M_{t}^{(i)}}$ using an efficient segmentation algorithm [11]. Each layer $\mathcal{L}_{t,m}^{(i)}$ corresponds to a continuous geometric surface patch with a coherent depth. Examples are shown in Fig. 3.
@@ 4.1.1 Tasks, datasets, and metrics
Following [72, 62], we evaluate on Sintel [2], Bonn [42], and KITTI [16]. Predicted depths are aligned to ground truth via scale-only alignment.
Camera pose estimation protocol. Following [72, 62], we compare on small-scale Sintel [2], ScanNet [8], and TUM RGB-D [52]. Predicted trajectories are aligned to ground truth via a Sim(3) transformation. For large-scale evaluation, we use KITTI Odometry [16] following [9].
Multi-view point map estimation protocol. We run evaluation on 7-Scenes [51] and NRGBD [67] with a keyframe sampling interval of 10, except for using an interval of 15 on NRGBD for [73, 24] due to their memory limits. Predicted point maps are registered to ground truth using the Umeyama algorithm (coarse Sim(3) alignment) followed by Iterative Closest Point (ICP) refinement.
Table 1: Video Depth Estimation on Sintel [2], Bonn [42] and KITTI [16]. We report Abs Rel and $\delta{<}1.25$.
@@
VGGT [59] 0.303 68.5 0.055 97.1 0.073 96.3
$\pi^{3}$ [64] 0.245 68.4 0.050 97.5 0.038 98.6
Spann3R [58] 0.622 42.6 0.144 81.3 0.198 73.7
CUT3R [62] 0.421 47.9 0.078 93.7 0.118 88.1
Point3R [66] 0.452 48.9 0.060 96.0 0.136 84.2
VGGT-SLAM [35] 0.424 56.0 0.076 93.2 0.136 81.8
StreamVGGT [73] 0.323 65.7 0.059 97.2 0.173 72.1
@@
Table 2: Camera Pose Estimation on Sintel [2], ScanNet [8], and TUM [52]. We report ATE, translational RPE, and rotational RPE.
@@
DUSt3R [63] 0.290 0.132 7.869 0.246 0.108 8.210 0.140 0.106 3.286
VGGT [59] 0.171 0.062 0.471 0.035 0.015 0.381 0.012 0.010 0.309
$\pi^{3}$ [64] 0.073 0.037 0.287 0.030 0.012 0.346 0.014 0.009 0.307
Spann3R [58] 0.329 0.110 4.471 0.096 0.023 0.661 0.056 0.021 0.591
CUT3R [62] 0.213 0.066 0.621 0.099 0.022 0.600 0.046 0.015 0.473
Point3R [66] 0.351 0.128 1.822 0.106 0.035 1.946 0.075 0.029 0.642
VGGT-SLAM [35] 0.303 0.128 6.883 0.070 0.049 2.447 0.030 0.020 1.567
StreamVGGT [73] 0.251 0.149 1.894 0.161 0.057 3.647 0.061 0.033 3.209
@@
Table 3: Large-scale Camera Pose Estimation on KITTI [16]. We report ATE (lower is better). CF: checkmark () indicates no calibration required; DR: checkmark () indicates dense reconstruction supported. We show metrics for sequence ID; Avg. is the mean across sequences. Seq. 01 corresponds to a high-speed driving sequence whose motion differs from others; Avg. reports the mean ATE excluding Seq. 01. OOM: CUDA out-of-memory, TL: tracking lost.
@@
ORB-SLAM2 (w/o LC) [38] 69.73 26.48 40.65 502.20 47.82 0.94 1.30 29.95 40.82 16.04 43.09 38.77 5.42
ORB-SLAM2 (w/ LC) [38] 54.82 9.46 6.03 508.34 14.76 1.02 1.57 4.04 11.16 2.19 38.85 8.39 6.63
LDSO [15] 22.43 23.50 9.32 11.68 31.98 2.85 1.22 5.10 13.55 2.96 129.02 21.64 17.36
DROID-VO [55] 54.19 51.19 98.43 84.20 108.80 2.58 0.93 59.27 64.40 24.20 64.55 71.80 16.91
DPVO [56] 53.61 57.70 113.21 12.69 123.40 2.09 0.68 58.96 54.78 19.26 115.90 75.10 13.63
DROID-SLAM [55] 100.28 75.85 92.10 344.60 107.61 2.38 1.00 118.50 62.47 21.78 161.60 72.32 118.70
DPV-SLAM [31] 53.03 57.19 112.80 11.50 123.53 2.50 0.81 57.80 54.86 18.77 110.49 76.66 13.65
DPV-SLAM++ [31] 25.75 27.14 8.30 11.86 39.64 2.50 0.78 5.74 11.60 1.52 110.90 76.70 13.70
VGGT [59] / / OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
$\pi^{3}$ [64] / / OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
MASt3R-SLAM [39] / / TL TL TL TL TL TL TL TL TL TL TL
CUT3R [62] / / OOM OOM OOM 148.07 22.31 OOM OOM OOM OOM OOM OOM
Fast3R [68] / / OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
VGGT-Long [9] 27.64 18.28 8.67 121.17 32.08 6.12 4.23 8.31 5.34 4.63 53.10 41.99 18.37
@@
Table 5: Ablation Studies for Layer-wise Scale Alignment (Sec. 3.2). wo// LSA denotes not using LSA component. w// SAM 2 denotes replacing the efficient segmentation algorithm [11] with a recent counterpart SAM 2. wo// $\mathcal{E}_{\mathrm{intra}}$ denotes removing the temporal propagation step from LSA.